# A sparsity test for multivariate Hawkes processes

### v1

@@ Abstract.
Multivariate Hawkes processes (mhp) are a class of point processes in which events at different coordinates interact through mutual excitation. The weighted adjacency matrix of the mhp encodes the strength of the relations, and shares its support with the causal graph of interactions of the process. We consider the problem of testing for causal relationships across the dimensions of a marked mhp. The null hypothesis is that a joint group of adjacency coefficients are null, corresponding to the absence of interactions. The alternative is that they are positive, and the associated interactions do exist. To this end, we introduce a novel estimation procedure in the context of a large sample of independent event sequences. We construct the associated likelihood ratio test and derive the asymptotic distribution of the test statistic as a mixture of $\chi^{2}$ laws. We offer two applications on financial datasets to illustrate the performance of our method. In the first, our test reveals a deviation from a static equilibrium in bidders’ strategies on retail online auctions. In the second, we uncover some factors at play in the dynamics of German intraday power prices.
@@ 2. Model and assumptions
Consider a rich enough probability space on which, for each $k\in[\![1,K]\!]$ , the event times $(t^{k}_{i})$ introduced above are defined as a sequence of increasing positive random variables and required to be distinct, meaning that jumps as different coordinates may not occur simultaneously. To each event time $t^{k}_{i}$ is associated a mark $(X^{k}_{i})$ with values in $\mathbb{R}^{m}$ and law $F^{k}(\mathop{}\!\mathrm{d}x)$. The multivariate counting process $(\boldsymbol{N}_{t})=(N_{1,t},\cdots,N_{K,t})$ for the $(t^{k}_{i})$ is then defined for any $t\geq 0$ and $k\in[\![1,K]\!]$ as
@@
and the counting measure $N_{k}(\mathop{}\!\mathrm{d}s,\mathop{}\!\mathrm{d}x)$ is the random Borel measure from (+)×$\mathcal{B}(\mathbb{R}^{+})\times\mathcal{B}(\mathbb{R})^{\otimes m}$ to $\mathbb{R}^{+}$
@@
Denote by $\mathbb{F}=(\mathcal{F}_{t})$ the natural filtration of $(\boldsymbol{N}_{t})$ and consider a $\mathbb{F}$-predictable coordinate-wise positive process $(\boldsymbol{\lambda}_{t})=(\lambda_{1,t},\cdots,\lambda_{K,t})$ From Jacod [27], under mild technical assumptions we may choose our probability space such that $(\boldsymbol{N}_{t})$ admits $(\boldsymbol{\Lambda}_{t})=(\int_{0}^{t}\boldsymbol{\lambda}_{s}\mathop{}\!\mathrm{d}s)$ as its compensator, that is, as the unique predictable increasing process such that $(\boldsymbol{N}_{t}-\boldsymbol{\Lambda}_{t})$ is a $\mathbb{F}$-local martingale. We refer to the comprehensive textbook of Daley & Vere-Jones [14] for more details on processes with stochastic intensity. This equips us with a characterisation of the (marked) mhp.
@@
$\boldsymbol{\lambda}_{t}=\mu\Big{(}\frac{t}{T}\Big{)}+\int_{[0,t)\times\mathbb{R}^{m}}\varphi(t-s,x)\boldsymbol{N}(\mathop{}\!\mathrm{d}s,\mathop{}\!\mathrm{d}x),$
@@
Define also $\mathop{}\!\mathrm{d}\boldsymbol{N}(s)=\boldsymbol{N}(\mathop{}\!\mathrm{d}s,\mathbb{R})$, and $\mathop{}\!\mathrm{d}\boldsymbol{M}^{(n)}(s)=\boldsymbol{M}(\mathop{}\!\mathrm{d}s,\mathbb{R})$. Then, from Definition 1,
@@
is a local martingale. While Jacod’s [27] theorem suffices to guarantee the existence of $(\boldsymbol{N}_{t})$, note that one can always explicitly construct the mhp such that the no simultaneous jump condition we required above is satisfied. See for instance Hawkes & Oakes [26] and Delattre et al. [15] for two equivalent representations, respectively as as cluster process and a Poisson sde solution. The nuance is important in that the jumps being distinct entails some useful simplification for the variation of the local martingale $(\boldsymbol{M}_{t})$, as transcribed in remark 1 below.
@@
We need more notation. Throughout, we denote by $\lVert\cdot\lVert_{p}$ the $\ell^{p}$-norm on $\mathbb{R}^{K}$ and by $\lVert\cdot\lVert_{L^{1}}$ the $L^{1}(\mathbb{R}^{+})$-norm on the positive integrable functions $\lVert\phi\lVert_{L^{1}}=\int_{0}^{\infty}\phi(s)\mathop{}\!\mathrm{d}s$. The set of squared real $K\times K$ matrices is labelled $\mathcal{M}_{K}(\mathbb{R})$ and for any $M\in\mathcal{M}_{K}(\mathbb{R})$, $\rho(M)$ is the spectral radius of $M$. For any open set $U\subset\mathbb{R}$, the set $\mathcal{C}^{0}(U)$ designates the continuous functions from $U$ to $\mathbb{R}$, and likewise $\mathcal{C}^{1}(U)$ designates the differentiable functions from $U$ to $\mathbb{R}$. Finally, for any $f=(f_{kl})\colon[0,\infty)\mapsto\mathcal{M}_{K}(\mathbb{R})$, we employ the notation
$F(\mathop{}\!\mathrm{d}x)=\text{Diag}(F^{j}(\mathop{}\!\mathrm{d}x))\hskip 14.22636pt\text{and}\hskip 14.22636pt\langle F,f\rangle=\int_{\mathbb{R}}f(\cdot,x)F(\mathop{}\!\mathrm{d}x),$
@@ 2.2. Estimation procedure and assumptions
The intensity is parameterised by $\theta$ over a compact subset $\Theta\subset\mathbb{R}^{p}$ with non empty interior. To simplify the expression for our test, we assume every parameter $\theta\in\Theta$ separates into $\theta=(\gamma,\alpha,\beta)$ where $\mu$ depends only on $\gamma\in\mathbb{R}^{K}$ and $\varphi$ expresses as
@@
Expression (8) and (9) are consistent with parameterisations found in the literature (see for instance Ogata [36] or Hardiman [25]), where $\Phi_{kl}$ is typically chosen (up to a re-normalisation) as the density of an exponential or power law with decay $\beta_{kl}$. The support of the matrix $\alpha$, that is, the matrix $h=(h_{kl})\in\mathcal{M}_{K}(\{0,1\})$ with $h_{kl}=0\iff\alpha_{kl}=0$, is then known as the adjacency of the mhp (see for instance Sulem [46]). Accordingly, $\alpha$ is known as its weighted adjacency (see Bacry et al. [5]). The null hypothesis conveniently re-phrases in terms of the matrix parameter $\alpha$ as
@@
    For any $\theta\in\Theta$, $k\in[\![1,p]\!]$ and $t\in[0,1]$, $\mu_{k}(t,\theta)>0.$
@@
    For any $\theta$ $\in\Theta$ and $k\in[\![1,p]\!]$, $\mu_{k}(\cdot,\theta)\in\mathcal{C}^{0}([0,1])$.
@@
Under Assumption 9, for every $t\in[0,T]$ and every $\theta\in\Theta$,
$\frac{1}{n}\mathbb{E}^{\theta}\left[\lambda^{(n)}(t,\theta)\right]=h(t,\theta):=\mu\Big{(}\frac{t}{T},\theta\Big{)}+\int_{0}^{t}\langle F,\Psi\rangle(t-s,\theta)\mu(s,\theta)\mathop{}\!\mathrm{d}s,$
where $\langle F,\Psi\rangle=\sum_{k\geq 1}\langle F,\varphi\rangle^{\star k}$
@@
    tales non singular values, where $h$ is the solution to the Volterra equation $h=\mu+\langle F,\varphi\rangle\star h$.
@@
for $(k,l)\in[\![1,K]\!]^{2}$. It suffices that $\alpha_{kl}=0$ in the true parameter to prevent $\beta_{kl}$ from being identified. Under the non degeneracy assumption 10, the size and structure of the decay parameter is hence constrained by those of the weighted adjacency. Note that this is a generic issue in parametric inference for mhps and not a specificity of our likelihood-based approach. We refer to Bonnet et al. [7] for a simple identifiability criterion in the context of exponential mhps. In general, we provide a consistent estimator for $I(\theta)$ in Lemma 11 on which assumption 10 can be numerically verified.
@@ 2.3. Likelihood-based inference
(16) $\mathcal{L}(n,\theta)=\sum_{k=1}^{K}\int_{0}^{T}\ln\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}N^{(n)}_{k}(s)-\int_{0}^{T}\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}s,$
where $(\boldsymbol{N}^{(n)}_{t})$ is the process defined by (6). We refer to Daley & Vere-Jones [14][Proposition 7.2.3 p. 232] for a proof of (16). In the rest of the paper, for any $\theta\in\Theta$, $\mathbb{P}(\theta)$ is the probability under which, for any $n\in\mathbb{N}^{\star}$, $(\boldsymbol{N}^{(n)}_{t})$ admits $(\lambda^{(n)}(t,\theta))$ as its intensity. The maximum likelihood estimator $\hat{\theta}_{n}$ (mle) for $\theta$ is then defined as the global argmax of the likelihood $\mathcal{L}(n,\cdot)$ over $\Theta$. For any sub-model $\Theta_{0}$, we define similarly $\hat{\theta}_{n}^{0}$ as the global argmax of $\mathcal{L}(n,\cdot)$ over $\Theta_{0}$. Note that $\hat{\theta}_{n}^{0}$ writes as a random vector of $\mathbb{R}^{d}$ with $p$ null coordinates. Our core problem is to derive the asymptotic distribution of the (log) likelihood ratio statistic (lrs)
@@ 3.2. Asymptotic of the likelihood ratio test
Let $\Theta_{0}$ be the sub-model (17) associated to a null hypothesis of the type (2). Without loss of generality we may assume that the null adjacencies correspond to the first $m$ parameter coordinates $\theta_{1},\cdots\theta_{m}$. Then, $\sqrt{n}(\Theta_{0}-\{\theta_{0}\})$ converges  [^1] In the sense of Painlevé-Kuratowski, see the introduction of Geyer [19]. to $H_{0}=\{0\}^{m}\times\mathbb{R}^{p-m}$ and $\sqrt{n}(\Theta-\{\theta_{0}\})$ to $H=(0,\infty)^{m}\times\mathbb{R}^{p-m}$. The asymptotic distribution of the lrs $\Lambda_{n}$ is then given as the difference between the squared distances of a certain Gaussian variable to $H_{0}$ and $H$.
@@
Under Assumptions 1 to 10, the likelihood ratio for the set of hypotheses (10) and (11) is asymptotically distributed as a chi-bar law:
@@ 6.1. Technical results
Let $\theta\in\Theta$. Under Assumption 1, for any $n\in\mathbb{N}^{\star}$, $\lambda^{(n)}_{k}$ is strictly positive and $\Phi(n^{-1}\lambda^{(n)}_{k,s}(\theta))$ is indeed well defined for any $k\in[\![1,K]\!]$. By the inverse triangular inequality, for any $t\in[0,T]$, and any $k\in[\![1,K]\!]$,
@@ 6.2. Proof of proposition 1
$-\sum_{k=1}^{K}\int_{0}^{T}\big{\{}\ln\Big{(}\frac{\frac{1}{n}\lambda^{(n)}_{k,s}(\theta_{0})}{\frac{1}{n}\lambda^{(n)}_{k,s}(\theta)}\Big{)}-\ln\Big{(}\frac{\bar{\lambda}(s,\theta_{0},\theta_{0})}{\bar{\lambda}(s,\theta,\theta_{0})}\Big{)}\big{\}}\frac{1}{n}\mathop{}\!\mathrm{d}N_{k,s}^{(n)}.$
@@ 7.1. Technical results
$\sum_{k=1}^{K}\int_{0}^{T}U\big{(}\frac{1}{n}Y^{(n)}_{k,ij}(s,\theta)^{(n)}\big{)}\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k,s}+\int_{0}^{T}V\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(s,\theta)^{(n)}\big{)}\mathop{}\!\mathrm{d}s,$
@@
$\frac{1}{\sqrt{n}}\boldsymbol{\mathcal{S}}(n,\theta_{0})$ $=\sum_{k=1}^{K}\int_{0}^{T}\frac{1}{\sqrt{n}}\frac{\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathop{}\!\mathrm{d}M^{(n)}_{k,s}$
(62) $=\sum_{k=1}^{K}\int_{0}^{T}U^{k}_{n}(s)\mathop{}\!\mathrm{d}M^{(n)}_{k,s},$
@@
$\big{(}\tilde{M}^{(n)}_{t}\big{)}=\big{(}\sum_{k=1}^{K}\int_{0}^{t}U^{k}_{n}(s)\mathop{}\!\mathrm{d}M^{(n)}_{k,s}\big{)}.$
@@
(63) $\left\langle\tilde{M}^{(n)},\tilde{M}^{(n)}\right\rangle(t)=\int_{0}^{t}\sum_{k=1}^{K}\frac{\left(\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)\right)^{\otimes 2}}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathop{}\!\mathrm{d}s.$
@@
$\tilde{M}^{(n)}_{\varepsilon}(t)=\sum_{k=1}^{K}\int_{0}^{t}U^{(n)}_{k}(s)\mathbb{1}_{\lvert U^{(n)}_{k}\lvert>\varepsilon}\mathop{}\!\mathrm{d}M^{(n)}_{k}(s),$
@@
$\langle\tilde{M}^{(n)}_{\varepsilon},\tilde{M}^{(n)}_{\varepsilon}\rangle(t)=\sum_{k=1}^{K}\int_{0}^{t}\frac{\left(\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)\right)^{\otimes 2}}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\sum_{l}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathbb{1}_{\lvert U^{(n)}_{k}\lvert>\varepsilon}\mathop{}\!\mathrm{d}s.$
@@
$\sup_{t\in[0,T]}U_{k}^{(n)}(t)=O_{\mathbb{P}(\theta_{0})}\big{(}\frac{1}{\sqrt{n}}\big{)}.$
@@
$\mathbb{P}\big{[}\lvert\sup_{t\in[0,T]}\lVert Z^{(n)}_{t}(\theta)\lVert_{1}-\sup_{t\in[0,T]}\lVert Z^{(n)}_{t}(\theta)\lVert_{1}\lvert>\eta d_{p^{*}}(\theta,\nu)\big{]}\leq 2K\exp\Big{[}-\frac{1}{2}\frac{\eta^{2}}{D^{2}}\Big{]}.$

### v2

@@ Abstract.
Multivariate Hawkes processes (mhp) are a class of point processes in which events at different coordinates interact through mutual excitation. The weighted adjacency matrix of the mhp encodes the strength of the relations, and shares its support with the causal graph of interactions of the process. We consider the problem of testing for causal relationships across the dimensions of a marked mhp. The null hypothesis is that a joint group of adjacency coefficients are null, corresponding to the absence of interactions. The alternative is that they are positive, and the associated interactions do exist. To this end, we introduce a novel estimation procedure in the context of a large sample of independent event sequences. We construct the associated likelihood ratio test and derive the asymptotic distribution of the test statistic as a mixture of $\chi^{2}$ laws. We offer two applications on financial datasets to illustrate the performance of our method. In the first one, our test reveals a deviation from a static equilibrium in bidders’ strategies on retail online auctions. In the second one, we uncover some factors at play in the dynamics of German intraday power prices.
@@ 2. Model and assumptions
Consider a rich enough probability space on which, for each $k\in[\![1,K]\!]$ , the event times $(t^{k}_{i})$ introduced above are defined as a sequence of increasing positive random variables and required to be distinct, meaning that jumps as different coordinates may not occur simultaneously. To each event time $t^{k}_{i}$ is associated a mark $(X^{k}_{i})$ with values in $\mathbb{R}^{m}$ and law $F^{k}(\mathop{}\!\mathrm{d}x)$. The $(X_{i}^{k})$ form an i.i.d collection of random variables. The multivariate counting process $(\boldsymbol{N}_{t})=(N_{1,t},\cdots,N_{K,t})$ for the $(t^{k}_{i})$ is then defined for any $t\geq 0$ and $k\in[\![1,K]\!]$ as
@@
and the counting measure $N_{k}(\mathop{}\!\mathrm{d}s,\mathop{}\!\mathrm{d}x)$ is the random Borel measure from $\mathcal{B}(\mathbb{R}^{+})\otimes\mathcal{B}(\mathbb{R})^{\otimes m}$ to $\mathbb{R}^{+}$
@@
Denote by $\mathbb{F}=(\mathcal{F}_{t})$ the natural filtration of $(\boldsymbol{N}_{t})$ and consider a $\mathbb{F}$-predictable coordinate-wise positive process $\lambda_{t}=(\lambda_{1,t},\cdots,\lambda_{K,t})$ From Jacod [27], under mild technical assumptions we may choose our probability space such that $(\boldsymbol{N}_{t})$ admits $(\boldsymbol{\Lambda}_{t})=(\int_{0}^{t}\lambda_{s}\mathop{}\!\mathrm{d}s)$ as its compensator, that is, as the unique predictable increasing process such that $(\boldsymbol{N}_{t}-\boldsymbol{\Lambda}_{t})$ is a $\mathbb{F}$-local martingale. We refer to the comprehensive textbook of Daley & Vere-Jones [14] for more details on processes with stochastic intensity. This equips us with a characterisation of the (marked) mhp.
@@
$\lambda_{t}=\mu\Big{(}\frac{t}{T}\Big{)}+\int_{[0,t)\times\mathbb{R}^{m}}\varphi(t-s,x)\boldsymbol{N}(\mathop{}\!\mathrm{d}s,\mathop{}\!\mathrm{d}x),$
@@
Define also $\mathop{}\!\mathrm{d}\boldsymbol{N}_{s}=\boldsymbol{N}(\mathop{}\!\mathrm{d}s,\mathbb{R}^{m})$, and $\mathop{}\!\mathrm{d}\boldsymbol{M}_{s}=\boldsymbol{M}(\mathop{}\!\mathrm{d}s,\mathbb{R}^{m})$. Then, from Definition 1,
@@
is a local martingale. While Jacod’s [27] theorem suffices to guarantee the existence of $(\boldsymbol{N}_{t})$, note that one can always explicitly construct the mhp such that the no simultaneous jump condition we required above is satisfied. See for instance Hawkes & Oakes [26] and Delattre et al. [15] for two equivalent representations, respectively as as cluster process and a Poisson sde solution. The jumps being distinct entails some useful simplification for the variation of the local martingale $(\boldsymbol{M}_{t})$, as transcribed in remark 1 below.
@@
We need more notation. Throughout, we denote by $\lVert\cdot\lVert_{p}$ the $\ell^{p}$-norm on $\mathbb{R}^{K}$ and by $\lVert\cdot\lVert_{L^{1}}$ the $L^{1}(\mathbb{R}^{+})$-norm on the positive integrable functions $\lVert\phi\lVert_{L^{1}}=\int_{0}^{\infty}\phi(s)\mathop{}\!\mathrm{d}s$. The set of squared real $K\times K$ matrices is labelled $\mathcal{M}_{K}(\mathbb{R})$ and for any $M\in\mathcal{M}_{K}(\mathbb{R})$, $\rho(M)$ is the spectral radius of $M$. For any open set $U\subset\mathbb{R}$, the set $\mathcal{C}^{0}(U)$ designates the continuous functions from $U$ to $\mathbb{R}$, and likewise $\mathcal{C}^{1}(U)$ designates the differentiable functions from $U$ to $\mathbb{R}$. Finally, for any $f=(f_{kl})\colon\mathbb{R}^{m}\mapsto\mathcal{M}_{K}(\mathbb{R}^{+})$, we employ the notation
$F(\mathop{}\!\mathrm{d}x)=\text{Diag}(F^{j}(\mathop{}\!\mathrm{d}x))\hskip 14.22636pt\text{and}\hskip 14.22636pt\langle F,f\rangle=\int_{\mathbb{R}^{m}}f(\cdot,x)F(\mathop{}\!\mathrm{d}x),$
@@ 2.2. Estimation procedure and assumptions
under the probability measure $\mathbb{P}(\theta)$ where $\theta$ lies in a compact subset $\Theta\subset\mathbb{R}^{d}$ with non empty interior. To simplify the expression for our test, we assume $\theta$ separates into $\theta=(\gamma,\alpha,\beta)$ where $\mu$ depends only on $\gamma\in\mathbb{R}^{K}$ and $\varphi$ expresses as
@@
Expression (8) and (9) are consistent with parameterisations found in the literature (see for instance Ogata [36] or Hardiman [25]). The support of the matrix $\alpha$, that is, the matrix $h=(h_{kl})\in\mathcal{M}_{K}(\{0,1\})$ with $h_{kl}=0\iff\alpha_{kl}=0$, is then known as the adjacency of the mhp (see for instance Sulem [46]). Accordingly, $\alpha$ is known as its weighted adjacency (see Bacry et al. [5]). The null hypothesis conveniently re-phrases in terms of the matrix parameter $\alpha$ as
@@
    For any $\theta\in\Theta$, $k\in[\![1,K]\!]$ and $t\in[0,1]$, $\mu_{k}(t,\theta)>0.$
@@
    For any $\theta$ $\in\Theta$ and $k\in[\![1,K]\!]$, $\mu_{k}(\cdot,\theta)\in\mathcal{C}^{0}([0,1])$.
@@
Under Assumption 9, for any $\theta_{0}\in\Theta$ and any $t\in[0,T]$
$\frac{1}{n}\mathbb{E}^{\theta_{0}}\big{[}\lambda^{(n)}(t,\theta_{0})\big{]}=h(t,\theta_{0}):=\mu\Big{(}\frac{t}{T},\theta_{0}\Big{)}+\int_{0}^{t}\langle F,\Psi\rangle(t-s,\theta_{0})\mu(s,\theta_{0})\mathop{}\!\mathrm{d}s,$
where $\mathbb{E}^{\theta}$ is the expectation under $\mathbb{P}(\theta)$ and $\langle F,\Psi\rangle=\sum_{k\geq 1}\langle F,\varphi\rangle^{\star k}$, with $\star k$ the $k$-fold convolution.
@@
    takes non singular values, where $h(t,\theta)=(h_{k}(t,\theta))$ is defined in Lemma 1.
@@
for $(k,l)\in[\![1,K]\!]^{2}$. It suffices that $\alpha_{kl}=0$ in the true parameter to prevent $\beta_{kl}$ from being identified. Under the non degeneracy assumption 10, the size and structure of the decay parameter are hence constrained by those of the weighted adjacency. We refer to Bonnet et al. [7] for a simple identifiability criterion in the context of exponential mhps. In general, we provide a consistent estimator for $I(\theta)$ in Lemma 11 on which assumption 10 can be numerically verified.
@@ 2.3. Likelihood-based inference
(16) $\mathcal{L}(n,\theta)=\sum_{k=1}^{K}\int_{0}^{T}\ln\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}N^{(n)}_{k,s}-\int_{0}^{T}\lambda^{(n)}_{k,s}(\theta)\mathop{}\!\mathrm{d}s,$
where $(\boldsymbol{N}^{(n)}_{t})$ is the process defined by (6). We refer to Daley & Vere-Jones [14][Proposition 7.2.3 p. 232] for a proof of (16). The maximum likelihood estimator $\hat{\theta}_{n}$ (mle) for $\theta$ is then defined as the global argmax of the likelihood $\mathcal{L}(n,\cdot)$ over $\Theta$. For any sub-model $\Theta_{0}$, we define similarly $\hat{\theta}_{n}^{0}$ as the global argmax of $\mathcal{L}(n,\cdot)$ over $\Theta_{0}$. Note that $\hat{\theta}_{n}^{0}$ writes as a random vector of $\mathbb{R}^{d}$ with $p$ null coordinates. Our core problem is to derive the asymptotic distribution of the (log) likelihood ratio statistic (lrs)
@@ 3.2. Asymptotic of the likelihood ratio test
Let $\Theta_{0}$ be the sub-model (17) associated to a null hypothesis of the type (2). Without loss of generality we may assume that the null adjacencies correspond to the first $p$ parameter coordinates $\theta_{1},\cdots\theta_{p}$. Then, $\sqrt{n}(\Theta_{0}-\{\theta_{0}\})$ converges  [^1] In the sense of Painlevé-Kuratowski, see the introduction of Geyer [19]. to $H_{0}=\{0\}^{p}\times\mathbb{R}^{d-p}$ and $\sqrt{n}(\Theta-\{\theta_{0}\})$ to $H=(0,\infty)^{p}\times\mathbb{R}^{d-p}$. The asymptotic distribution of the lrs $\Lambda_{n}$ is then given as the difference between the squared distances of a certain Gaussian variable to $H_{0}$ and $H$.
@@
Under Assumptions 1 to 10, the likelihood ratio for the set of hypotheses (10) and (11) is asymptotically distributed under any $\mathbb{P}(\theta_{0})$ with $\theta_{0}\in\Theta_{0}$ as a chi-bar law:
@@ 6.1. Technical results
Let $\theta\in\Theta$. Under Assumption 1, for any $n\in\mathbb{N}^{\star}$, $\lambda^{(n)}_{k}$ is strictly positive and $\Phi(n^{-1}\lambda^{(n)}_{k,s}(\theta))$ is indeed well defined for any $k\in[\![1,K]\!]$. By the inverse triangle inequality, for any $t\in[0,T]$, and any $k\in[\![1,K]\!]$,
@@ 6.2. Proof of proposition 1
$-\sum_{k=1}^{K}\int_{0}^{T}\big{\{}\ln\Big{(}\frac{\frac{1}{n}\lambda^{(n)}_{k,s}(\theta_{0})}{\frac{1}{n}\lambda^{(n)}_{k,s}(\theta)}\Big{)}-\ln\Big{(}\frac{\bar{\lambda}_{k}(s,\theta_{0},\theta_{0})}{\bar{\lambda}_{k}(s,\theta,\theta_{0})}\Big{)}\big{\}}\frac{1}{n}\mathop{}\!\mathrm{d}N_{k,s}^{(n)}.$
@@ 7.1. Technical results
$\sum_{k=1}^{K}\int_{0}^{T}U\big{(}\frac{1}{n}Y^{(n)}_{k,ij}(s,\theta)\big{)}\frac{1}{n}\mathop{}\!\mathrm{d}N^{(n)}_{k,s}+\int_{0}^{T}V\big{(}\frac{1}{n}Y_{k,ij}^{(n)}(s,\theta)\big{)}\mathop{}\!\mathrm{d}s,$
@@
$\frac{1}{\sqrt{n}}\boldsymbol{\mathcal{S}}(n,\theta_{0})$ $=\sum_{k=1}^{K}\int_{0}^{T}\frac{1}{\sqrt{n}}\frac{\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,x,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\varphi_{kl}(s-u,x,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathop{}\!\mathrm{d}M^{(n)}_{k,s}$
(62) $=\sum_{k=1}^{K}\int_{0}^{T}U^{(n)}_{k}(s)\mathop{}\!\mathrm{d}M^{(n)}_{k,s},$
@@
$\big{(}\tilde{M}^{(n)}_{t}\big{)}=\big{(}\sum_{k=1}^{K}\int_{0}^{t}U^{(n)}_{k}(s)\mathop{}\!\mathrm{d}M^{(n)}_{k,s}\big{)}.$
@@
(63) $\left\langle\tilde{M}^{(n)},\tilde{M}^{(n)}\right\rangle(t)=\int_{0}^{t}\sum_{k=1}^{K}\frac{\left(\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)\right)^{\otimes 2}}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\varphi_{kl}(s-u,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathop{}\!\mathrm{d}s.$
@@
$\tilde{M}^{(n)}_{\varepsilon}(t)=\sum_{k=1}^{K}\int_{0}^{t}U^{(n)}_{k}(s)\mathbb{1}_{\lvert U^{(n)}_{k}\lvert>\varepsilon}\mathop{}\!\mathrm{d}M^{(n)}_{k,s},$
@@
$\langle\tilde{M}^{(n)}_{\varepsilon},\tilde{M}^{(n)}_{\varepsilon}\rangle(t)=\sum_{k=1}^{K}\int_{0}^{t}\frac{\left(\partial_{\theta}\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\partial_{\theta}\varphi_{kl}(s-u,x,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)\right)^{\otimes 2}}{\mu_{k}(s,\theta_{0})+\int_{0}^{s}\int_{\mathbb{R}^{m}}\sum_{l}\varphi_{kl}(s-u,x,\theta_{0})\frac{1}{n}N^{(n)}_{l}(\mathop{}\!\mathrm{d}u,\mathop{}\!\mathrm{d}x)}\mathbb{1}_{\lvert U^{(n)}_{k}\lvert>\varepsilon}\mathop{}\!\mathrm{d}s.$
@@
$\sup_{t\in[0,T]}\lvert U_{k}^{(n)}(t)\rvert=O_{\mathbb{P}(\theta_{0})}\big{(}\frac{1}{\sqrt{n}}\big{)}.$
@@
$\mathbb{P}\big{[}\lvert\sup_{t\in[0,T]}\lVert Z^{(n)}_{t}(\theta)\lVert_{1}-\sup_{t\in[0,T]}\lVert Z^{(n)}_{t}(\nu)\lVert_{1}\lvert>\eta d_{p^{*}}(\theta,\nu)\big{]}\leq 2K\exp\Big{[}-\frac{1}{2}\frac{\eta^{2}}{D^{2}}\Big{]}.$